What is the x-intercept of the equation 4x-8y=8?
A.(8,0)
B.(1/2,0)
C.(-2,0)
D.(2,0)
(subject is intercepts & zeros, but it wasn't available)

Answer by stanbon(75887)

What is the x-intercept of the equation 4x-8y=8?
A.(8,0)
B.(1/2,0)
C.(-2,0)
D.(2,0)
---------------------
To find the x-intercept, let y = 0; then solve for "x":
4x-8y = 8
4x-0 = 8
4x = 8
x = 2
That is the x-intercept.
============================
Cheers,
Stan H.